using ϵ−δ definition, show that

lim x→2 (1/2 x2 - x + 1)


1
Expert's answer
2022-03-19T02:39:41-0400

Let "\varepsilon >0" be fixed. We may assume that "\varepsilon \leqslant \frac{1}{2}" , otherwise we take "\delta" for "\varepsilon =\frac{1}{2}" . Solve

"\left| \frac{1}{2}x^2-x+1-1 \right|<\varepsilon \Leftrightarrow \left| x^2-2x \right|<2\varepsilon \Leftrightarrow \\\Leftrightarrow \left\{ \begin{array}{c} x^2-2x-2\varepsilon <0\\ x^2-2\delta δ x+2\varepsilon >0\\\end{array} \right. \Leftrightarrow \left\{ \begin{array}{c} x\in \left( 1-\sqrt{1+2\varepsilon},1+\sqrt{1+2\varepsilon} \right)\\ x\in \left( -\infty ,1-\sqrt{1-2\varepsilon} \right) \cup \left( 1+\sqrt{1-2\varepsilon},+\infty \right)\\\end{array} \right."

Thus

"x\in \left( 1+\sqrt{1-2\varepsilon},1+\sqrt{1+2\varepsilon} \right) \Rightarrow \left| \frac{1}{2}x^2-x+1-1 \right|<\varepsilon"

Let "\delta =\min \left( \sqrt{1+2\varepsilon}-1,1-\sqrt{1-2\varepsilon} \right) >0"

Then "x\in \left( 2-\delta ,2+\delta \right) \Rightarrow \left| \frac{1}{2}x^2-x+1-1 \right|<\varepsilon"

which by the definition proves

"\underset{x\rightarrow 2}{\lim}\left( \frac{1}{2}x^2-x+1 \right) =1"